# Localized Band Gap Resonances in Defect-embedded Monoatomic Chains

### v1

@@ III.2 Customized Acoustic absorber
Figure 6: Customized Acoustic absorber. (a) An edge DMoA-based AMM design with an engineered edge defect mode at $\Omega_{\mathrm{d}}=1$. (b) Velocity response, (c) the corresponding FFTs, and (d) Energy distribution at mass indices – (i) $j=1$, (ii) $j=(\mathrm{N}+1)/2$, and (iii) $j=\mathrm{N}$ ($\mathrm{N}=51$), of three different AMM configurations based on the base MoA (black), and two edge DMoAs with defect properties $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}\rightarrow$$\{1,0.382\}$ ($\square$), and $\{2.618,2\}$ ($\bullet$) used in Fig. 3.
@@
Here we present an edge DMoA-based AMM that interacts with a broadband acoustic field, $\mathrm{F}(\bar{t})$, and manifests a prescribed narrow-band response at the edge defect while simultaneously customizing the transmission of PB frequencies through the AMM bulk. Fig. 6a shows the mass-spring configuration of an edge DMoA-based AMM subject to an external force, $\mathrm{F}(\bar{t})$ at the edge defect location. $\mathrm{F}(\bar{t})$ is chosen as a representative broadband acoustic signal (as illustrated by the green FFT curve in Fig. 6(c-i)) which can originate from a variety of environmental sources, e.g., vehicular traffic, fluid flow. A customized acoustic interaction of the AMM with this signal can be useful in mitigating or harvesting this ambient energy. To investigate this capability, we numerically simulate the response of the base MoA (black) devoid of material defects and two AMM configurations described by edge DMoA parameters chosen from Fig. 3 (blue and red). Fig. 6(b-d) plot, respectively, the velocity response of the three AMM configurations, the FFTs of these velocity signals (normalized by peak the largest FFT amplitude of all velocity signals), and the energy distribution at mass indices – (i) $j=1$ (edge defect), (ii) $j=(\mathrm{N}+1)/2$ (midpoint), and (iii) $j=\mathrm{N}$ (last mass).
@@
The velocity responses and the corresponding FFTs of the base MoA in Fig. 6(b,c) indicate its behavior as a passive acoustic filter that allows propagation of frequency components of $\mathrm{F}(\bar{t})$ in the PB range, $\Omega\in[\sqrt{k_{\mathrm{g,r}}},\sqrt{4+k_{\mathrm{g,r}}}]$ while attenuating frequencies in the BGs. Alternately, both the AMMs manifest a narrow-band velocity response, centered at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ at $j=1$. Though PB frequency signatures of $\mathrm{F}(\bar{t})$ persist at $j=(\mathrm{N}+1)/2,\mathrm{N}$ in the AMMs, the level of transmission is significantly lower compared to the base MoA. To analyze this contrasting behavior, we quantify the total energy at a given mass index:
@@
Consistent with the velocity responses in Fig. 6(b-ii,iii), the total energy bars, $\mathrm{TE}_{(\mathrm{N}+1)/2,\mathrm{N}}$ in Fig. 6(d-ii,iii), are significantly lower in both the AMMs compared to the base MoA indicating a reduction in transmission into the material bulk. This contrast in the transmission is facilitated by a superior energy localization at the edge defect in both AMMs, as evident from the higher energy content in Fig. 6(d-i). Furthermore, each AMM manifests a distinct energy localization behavior commensurate with the prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ scales from Fig. 3. The persistence of these prescribed velocity scales throughout the AMMs (Fig.6(b,d-ii,iii)) reiterates the importance of the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ as a critical behavioral parameter that customizes both the localized response at the defect location and the energy transmission into the bulk.
Additional modifications to the AMM, e.g., adding more defects[26], secondary branching of the mass-spring chain[49], can be made according to operational requirements to further customize the AMM behavior without affecting the characteristics of the existing edge defect under reasonable proximity limits(II.3). Therefore, our modified analytical method offers a systematic strategy to design multi-functional defect-embedded AMMs based on simple mass-spring reduced-order models.
IV Conclusion
In summary, we present a modified analytical approach that allows us to engineer BG resonances with prescribed characteristics in AMMs by embedding material defects and breaking the periodicity of the underlying material matrix. The perturbed Toeplitz matrix eigenvalue problem informs our analytical approach which accurately quantifies the acoustic characteristics of single and multiple defect-embedded AMMs with reasonable proximity limits. Furthermore, we demonstrate the benefits of defect modes through two DMoA-based AMM examples that feature an in-house virtual ground, and a customized acoustic absorption of an incident broadband acoustic field. The proposed strategy can be readily extended to lattices operating under general boundary conditions[28, 30] and polyatomic unit-cell[28, 29] based lattices as required, and augments the toolkit for designing defect-based AMMs.
@@ Acknowledgments
VR and KM acknowledge funding from AFOSR MURI under grant number FA9550-23-1-0299.
@@
The matrix equation of motion for an edge defect-embedded grounded monoatomic chain (edge DMoA), with the temporal frequency normalized by $\omega_{0}=\sqrt{k/m}$, is given as:
@@
Assuming harmonic motion, $\mathbf{V}=\hat{\mathbf{V}}e^{i\Omega\bar{t}}$, the eigenvalue problem for the DMoA can be formulated as $|\mathbf{K}_{\mathrm{d}}|=\mathbf{0}$, where,
@@
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$ (A3)
@@
$U_{n}(x)=\frac{\sin(2(n+1)\kappa)}{\sin(2\kappa)},$ (A5)
@@
In the above equations, $\kappa$, represents the spatial frequency (non-dimensional wavenumber) associated with a given harmonic temporal frequency, $\Omega$. Eq. (A4) is an implicit equation that cannot be solved to obtain an explicit analytical form relating the DMoA material parameters to $\kappa$ and $\Omega$. Hence, we rearrange the LHS in Eq. (A4) to generate the modified equation, Eq. (3) listed in the main text.
@@
Eq. (3) is solved numerically by sweeping through a given range of frequencies, $\Omega$ to identify points of intersections between functions $f$ and $g$. A graphical representation of the same is shown in Fig. A1. Despite the U-transformation method utilized by Cai et al.[32] for the case of a bulk defect and the perturbed Toeplitz matrix-based method that we propose for the edge defect being fundamentally different solution approaches, a noteworthy similarity between the procedures is that they both involve implicit equations in the final step that need to be solved numerically to obtain the defect eigenfrequencies and eigenmode shapes.
@@
Though, the design space of the defect properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ probed in Fig. 2 in the main text manifest a single defect mode, other combinations, especially cases where $m_{\mathrm{r}}<1$ and $k_{\mathrm{gd,r}}<k_{\mathrm{g,r}}$ can manifest an additional defect mode in $2^{\mathrm{nd}}$ BG of the MoA, localized at the same edge defect location as the existing defect mode. Fig. A2 presents the (a) defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and (b) the decay rates, $\kappa_{I}$ when probing the defect material properties, $m_{\mathrm{r}}\in[0.04,1]$ and $k_{\mathrm{gd,r}}=0$. Not only do the analytical and numerical values of the defect frequencies and decay rates agree well, but we also notice that below the value, $m_{\mathrm{r}}\approx 0.33$ the eigenfrequencies of the finite DMoA undergo a bifurcation, where the highest PB frequency relocates into the $2\mathrm{nd}$ BG, appearing as a second defect resonance frequency in the DMoA chain.
Figure A2: Transition of a DMoA with material defect at a single boundary from manifesting a single defect mode to two defect modes. (a) Defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and, the corresponding (b) decay rates, $\kappa_{I}$. ($k_{\mathrm{gd,r}}=0$)
@@
The matrix equation governing the forced response of a DMoA subject to an external, single-frequency harmonic forcing is given as:
@@
where, $\mathbf{F}_{0}$ represents an $\mathrm{N}\times 1$ column vector with complex force excitation amplitudes (i.e., $\mathbf{F}_{0,j}\in\mathbb{C}$ for $j=1,2,\cdots,\mathrm{N}$), $\Omega_{\mathrm{exc}}$ represents the non-dimensional excitation frequency, and $\mathrm{c.c.}$ stands for the complex conjugate terms. If the excitation frequency coincides with one of the natural frequencies of the systems, i.e., $\Omega_{\mathrm{exc}}=\Omega_{j}$, the general solution ansatz for the displacement of the DMoA at resonance is of the form:
@@
Now, assume an external sinusoidal forcing, $\mathbf{F}=[\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t}),0,\cdots,0]^{\mathrm{T}}$, applied on the DMoA boundary housing the (edge) defect. The complex forcing amplitudes are expressed as:
@@
Substituting, the above equation into Eq. (A10b)a, we get $b=0$, as $\Omega_{\mathrm{d}}$ and $\hat{\mathbf{V}}_{\mathrm{d}}$ are a solution to the eigenvalue problem $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}_{\mathrm{d}}=0$. Therefore, the DMoA response to the given sinusoidal forcing can be written as:
@@
A4 Multiple defects
@@
Figure A3: Multi-DMoA chain with defect frequencies in $2^{\mathrm{nd}}$ BG of the MoA. (a) Absolute percentage error contours for a Multi-DMoA ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$, (b) Eigenmode shapes of the Multi-DMoA with one edge ($x_{\mathrm{d1}}=0$) and one bulk ($x_{\mathrm{d2}}$) defect. Multi-DMoA material parameters - $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$, $k_{\mathrm{r}}=1$ and $\{m_{r},k_{\mathrm{gd,r}}\}=\{1,6\}$
@@
The modified analytical approach presented in this article can be applied to determine the defect resonance frequencies of a Multi-DMoA provided the additional defects remain relatively dormant at the resonance frequency of a considered defect. To verify the accuracy of the analytical estimate of $\Omega_{\mathrm{d}}$, Fig. 4 in the main text plots the absolute percentage error b.w. the numerical and analytical values of the defect frequency given by,
@@
When the defect frequencies lie in the $1^{\mathrm{st}}$ BG of the MoA, we notice a small surge in the $\Omega_{\mathrm{d1}}$ error when at least one of the penultimate masses are prescribed defects, i.e, $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$. This error surge is attributed to the spatial in-phase vibration associated with the defect modes in the $1^{\mathrm{st}}$ BG and the significant impact of defect proximity to the finite boundary. Alternately, Fig. A3 explores a Multi-DMoA chain with defect modes now in the $2^{\mathrm{nd}}$ BG of a MoA chain. The error contours at $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$ in Fig. A3a are virtually indistinguishable from the surrounding configurations (blue region - $0.003\%$, surge region - $0.29\%$) as opposed to the error contours in Fig. 4a (blue region - $0.004\%$, surge region - $3.9\%$). The reduction in the error surge is indeed a result of improved accuracy of the analytical model and out-of-phase vibration of the defect and boundary masses as the absolute error, $\Delta\Omega_{\mathrm{d}}$ also decreases for defect modes in the $2^{\mathrm{nd}}$ BG.
A5 AMM with a Virtual Ground
The modified stiffness matrix,

### v2

@@ 3.2 Customized acoustic absorber
Figure 7: Customized acoustic absorber. (a) An edge DMoA-based PnC design with an engineered edge defect mode at $\Omega_{\mathrm{d}}=1$. (b) Velocity response, (c) the corresponding FFTs, and (d) Energy distribution at mass indices – (i) $j=1$, (ii) $j=(\mathrm{N}+1)/2$, and (iii) $j=\mathrm{N}$ ($\mathrm{N}=51$), of three different PnC configurations - the base MoA PnC (black), and two edge DMoA PnCs with defect properties $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}\rightarrow$$\{1,0.382\}$ ($\square$), and $\{2.618,2\}$ ($\bullet$) used in Fig. 3.
@@
Here we present an edge DMoA-based PnC that interacts with a broadband acoustic field, $\mathrm{F}(\bar{t})$, and manifests a prescribed narrow-band response at the edge defect while simultaneously customizing the transmission of PB frequencies through the PnC bulk. Fig. 7a shows the mass-spring configuration of an edge DMoA-based PnC subject to an external force, $\mathrm{F}(\bar{t})$ at the edge defect location. $\mathrm{F}(\bar{t})$ is chosen as a representative broadband acoustic signal (as illustrated by the green FFT curve in Fig. 7(c-i)) which can originate from a variety of environmental sources, e.g., vehicular traffic, machine noise in factories, fluid flow. A customized acoustic interaction of the PnC with this signal can be useful in mitigating or harvesting this ambient energy. To investigate this capability, we numerically simulate the response of the base MoA PnC (black) devoid of material defects and two DMoA PnC configurations described by edge defect parameters chosen from Fig. 3 (blue and red). Fig. 7(b-d) plot, respectively, the velocity response of the three PnC configurations, the FFTs of these velocity signals (normalized by peak the largest FFT amplitude of all velocity signals), and the energy distribution at mass indices – (i) $j=1$ (edge defect), (ii) $j=(\mathrm{N}+1)/2$ (midpoint), and (iii) $j=\mathrm{N}$ (last mass).
@@
The velocity responses and the corresponding FFTs of the MoA PnC in Fig. 7(b,c) indicate its behavior as a passive acoustic filter that allows propagation of frequency components of $\mathrm{F}(\bar{t})$ in the PB range, $\Omega\in[\sqrt{k_{\mathrm{g,r}}},\sqrt{4+k_{\mathrm{g,r}}}]$ while attenuating frequencies in the BGs. Alternately, both the PnCs manifest a narrow-band velocity response, centered at the prescribed defect frequency, $\Omega_{\mathrm{d}}=1$ at $j=1$. Though PB frequency signatures of $\mathrm{F}(\bar{t})$ persist at $j=(\mathrm{N}+1)/2,\mathrm{N}$ in the PnCs, the level of transmission is significantly lower compared to the MoA PnC. To analyze this contrasting behavior, we quantify the total energy at a given mass index:
@@
Consistent with the velocity responses in Fig. 7(b-ii,iii), the total energy bars, $\mathrm{TE}_{(\mathrm{N}+1)/2,\mathrm{N}}$ in Fig. 7(d-ii,iii), are significantly lower in both the PnCs compared to the MoA PnC indicating a reduction in transmission into the material bulk. This contrast in the transmission is facilitated by a superior energy localization at the edge defect in both PnCs, as evident from the higher energy content in Fig. 7(d-i). Furthermore, each PnC manifests a distinct energy localization behavior commensurate with the prescribed $\dot{\mathrm{V}}_{\mathrm{E}}$ scales from Fig. 3. The persistence of these prescribed velocity scales throughout the PnCs (Fig.7(b,d-ii,iii)) reiterates the importance of the velocity amplitude envelope, $\dot{\mathrm{V}}_{\mathrm{E}}$ as a critical behavioral parameter that customizes both the localized response at the defect location and the energy transmission into the bulk.
Additional modifications to the PnC, e.g., adding more defects[26], secondary branching of the mass-spring PnC[49], can be made according to operational requirements to further customize the PnC behavior without affecting the characteristics of the existing edge defect under reasonable proximity limits(2.4). Therefore, our modified analytical method offers a systematic strategy to design multi-functional defect-embedded PnCs based on simple mass-spring reduced-order models.
4 Conclusion
In summary, we present a modified analytical approach that allows us to engineer BG resonances with prescribed characteristics in PnCs by embedding material defects and breaking the periodicity of the underlying homogeneous material. The perturbed Toeplitz matrix eigenvalue problem informs our analytical approach which accurately quantifies the acoustic characteristics of 1D and 2D DMoA PnCs with single and multiple defect-embedded with reasonable proximity limits. Furthermore, we demonstrate the benefits of defect modes through two DMoA-based PnC examples that feature an in-house virtual ground, and a customized acoustic absorption of an incident broadband acoustic field. The proposed strategy can be readily extended to lattices operating under general boundary conditions[28, 30] and polyatomic unit-cell[29] based lattices, and study topologically protected modes[29], given at least a diatomic PnC. Therefore, the proposed method augments the toolkit for designing defect-based PnCs.
@@ Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0299 and award number FA9550-21-1-0182.
@@
The matrix equation of motion for an edge defect-embedded grounded monoatomic phononic crystal (edge DMoA PnC), with the temporal frequency normalized by $\omega_{0}=\sqrt{k/m}$, is given as:
@@
Assuming harmonic motion, $\mathbf{V}=\hat{\mathbf{V}}e^{i\Omega\bar{t}}$, the eigenvalue problem for the DMoA PnC can be formulated as $|\mathbf{K}_{\mathrm{d}}|=\mathbf{0}$, where,
@@
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$
@@
$U_{n}(x)=\frac{\sin(2(n+1)\kappa)}{\sin(2\kappa)},$
@@
In the above equations, $\kappa$, represents the spatial frequency (non-dimensional wavenumber) associated with a given harmonic temporal frequency, $\Omega$. Eq. (A3) is an implicit equation that cannot be solved to obtain an explicit analytical form relating the DMoA PnC material parameters to $\kappa$ and $\Omega$. Hence, we rearrange the LHS in Eq. (A3) to generate the modified equation, Eq. (3) listed in the main text.
@@
Eq. (3) is solved numerically by sweeping through a given range of frequencies, $\Omega$ to identify points of intersections between functions $f$ and $g$. A graphical representation is shown in Fig. A1. Despite the U-transformation method utilized by Cai et al.[32] for the case of a bulk defect and the perturbed Toeplitz matrix-based method that we propose for the edge defect being fundamentally different solution approaches, a noteworthy similarity between the procedures is that they both involve implicit equations in the final step that need to be solved numerically to obtain the defect eigenfrequencies and eigenmode shapes.
Damped DMoA PnCs
The normalized matrix equation of motion for a damped DMoA PnC is given as:
$\mathbf{M}\ddot{\mathbf{V}}+\mathbf{C}\dot{\mathbf{V}}+\mathbf{K}\mathbf{V}=\mathbf{0}$
where, the general damping matrix has the form,
$\mathbf{C}=c_{\mathrm{on,r}}\mathbf{I}+c_{\mathrm{in,r}}\mathbf{K}_{\mathrm{in}}=\begin{bmatrix}c_{\mathrm{on,r}}+c_{\mathrm{in,r}}&-c_{\mathrm{in,r}}&0&\dots&0\\ -c_{\mathrm{in,r}}&c_{\mathrm{on,r}}+2c_{\mathrm{in,r}}&-c_{\mathrm{in,r}}&\dots&0\\ 0&-c_{\mathrm{in,r}}&\ddots&\ddots&\vdots\\ \vdots&\vdots&\vdots&\ddots&-c_{\mathrm{in,r}}\\ 0&0&\dots&-c_{\mathrm{in,r}}&c_{\mathrm{on,r}}+c_{\mathrm{in,r}}\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},$ (A4)
$\mathbf{I}=\begin{bmatrix}1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},\mathbf{K}_{\mathrm{in}}=\begin{bmatrix}1&-1&0&\cdots&0\\ -1&2&-1&\cdots&0\\ 0&-1&&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&-1\\ 0&0&\cdots&-1&1\end{bmatrix}_{\mathrm{N}\times\mathrm{N}},$
Consequently, the dynamic stiffness matrix and matrix coefficients can be modified as - $\mathbf{K}_{\mathrm{d}}=\mathbf{K}+i\Omega\mathbf{C}-\Omega^{2}\mathbf{M}$
$d=2+k_{\mathrm{g,r}}+i\Omega(c_{\mathrm{on,r}}+2c_{\mathrm{in,r}})-\Omega^{2},a=-(1+i\Omega c_{\mathrm{in,r}}),\beta=-(1+i\Omega c_{\mathrm{in,r}}),$
$\alpha=(k_{\mathrm{gd,r}}-k_{\mathrm{g,r}})-(1+i\Omega c_{\mathrm{in,r}})+(1-m_{\mathrm{r}})\Omega^{2}$
@@
Though, the design space of the defect properties, $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$ probed in Fig. 2 in the main text manifest a single defect mode, other combinations, especially cases where $m_{\mathrm{r}}<1$ and $k_{\mathrm{gd,r}}<k_{\mathrm{g,r}}$ can manifest an additional defect mode in $2^{\mathrm{nd}}$ BG of the MoA PnC, localized at the same edge defect location as the existing defect mode. Fig. A2 presents the (a) defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and (b) the decay rates, $\kappa_{I}$ when probing the defect material properties, $m_{\mathrm{r}}\in[0.04,1]$ and $k_{\mathrm{gd,r}}=0$. Not only do the analytical and numerical values of the defect frequencies and decay rates agree well, but we also notice that below the value, $m_{\mathrm{r}}\approx 0.33$ the eigenfrequencies of the finite DMoA PnC undergo a bifurcation, where the highest PB frequency relocates into the $2\mathrm{nd}$ BG, appearing as a second defect resonance frequency in the DMoA PnC.
Figure A2: Transition of a DMoA PnC with material defect at a single boundary from manifesting a single defect mode to two defect modes. (a) Defect frequencies, $\Omega_{\mathrm{d1,d2}}$ and, the corresponding (b) decay rates, $\kappa_{I}$. ($k_{\mathrm{gd,r}}=0$)
@@
The matrix equation governing the forced response of a DMoA PnC subject to an external, single-frequency harmonic forcing is given as:
@@
where, $\mathbf{F}_{0}$ represents an $\mathrm{N}\times 1$ column vector with complex force excitation amplitudes (i.e., $\mathbf{F}_{0,j}\in\mathbb{C}$ for $j=1,2,\cdots,\mathrm{N}$), $\Omega_{\mathrm{exc}}$ represents the non-dimensional excitation frequency, and $\mathrm{c.c.}$ stands for the complex conjugate terms. If the excitation frequency coincides with one of the natural frequencies of the systems, i.e., $\Omega_{\mathrm{exc}}=\Omega_{j}$, the general solution ansatz for the displacement of the DMoA PnC at resonance is of the form:
@@
Now, assume an external sinusoidal forcing, $\mathbf{F}=[\mathrm{F}_{0}\sin(\Omega_{\mathrm{d}}\bar{t}),0,\cdots,0]^{\mathrm{T}}$, applied on the DMoA PnC boundary housing the (edge) defect. The complex forcing amplitudes are expressed as:
@@
Substituting, the above equation into Eq. (A9b)a, we get $b=0$, as $\Omega_{\mathrm{d}}$ and $\hat{\mathbf{V}}_{\mathrm{d}}$ are a solution to the eigenvalue problem $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}_{\mathrm{d}}=0$. Therefore, the DMoA PnC response to the given sinusoidal forcing can be written as:
@@
A4 Steady-State Velocity Amplitude
The matrix equation governing the forced response of a damped DMoA PnC can be written as:
$\mathbf{M}\ddot{\mathbf{V}}+\mathbf{C}\dot{\mathbf{V}}+\mathbf{K}\mathbf{V}=\frac{\mathbf{F}_{0}}{2}e^{i\Omega_{\mathrm{exc}}\bar{t}}+\mathrm{c.c.},$ (A12)
The general ansatz for the displacement of the DMoA PnC is of the form:
$\mathbf{V}(\bar{t})=\sum_{j=1}^{N}\left(\frac{A_{j}}{2}\right)e^{(-\xi_{j}\Omega_{n,j}+i\Omega_{j})\bar{t}}\hat{\mathbf{V}}_{j}+\left(\frac{B_{\mathrm{exc}}}{2}\right)e^{i\Omega_{j}\bar{t}}\hat{\mathbf{V}}_{\mathrm{exc}}+\mathrm{c.c.},$ (A13)
where, $\xi_{j}$ is the modal damping ratio, $\Omega_{j}$ is the $j^{\mathrm{th}}$ damped natural frequency, $\Omega_{n,j}=\left|\Omega_{j}/\sqrt{1-\xi_{j}^{2}}\right|$, and coefficients $A_{j},B\in\mathbb{C}$. When exciting at the defect resonance frequency, $\Omega_{j}=\Omega_{\mathrm{d}}$ the coefficients $A_{j}=0\ \forall\ j\neq\mathrm{d}$. Applying initial displacement and velocity conditions - $\mathbf{V}(0)=\mathbf{0}$ and $\dot{\mathbf{V}}(0)=\mathbf{0}$, assuming, $A_{\mathrm{d}}=a_{R}+ia_{I}$ and $B_{\mathrm{exc}}=B_{\mathrm{d}}=b_{R}+ib_{I}$:
$\mathbf{V}(0)=\mathbf{0}\Longrightarrow$ $\mathrm{Re}(A_{\mathrm{d}}+B_{\mathrm{d}})=0$
$\Longrightarrow$ $a_{R}=-b_{R},\mathrm{for}\ a_{R}\in\mathbb{R},$ (A14a)
$\dot{\mathbf{V}}(0)=\mathbf{0}\Longrightarrow$ $\xi_{\mathrm{d}}\Omega_{n,\mathrm{d}}\mathrm{Re}(A_{\mathrm{d}})=\Omega_{\mathrm{d}}\mathrm{Im}(A_{\mathrm{d}}+B_{\mathrm{d}})$
$\Longrightarrow$ $\xi_{\mathrm{d}}\Omega_{n,\mathrm{d}}a_{R}=\Omega_{\mathrm{d}}(a_{I}+b_{I}),\mathrm{for}\ a_{I},b_{I}\in\mathbb{R},$ (A14b)
Substituting the ansatz - $\mathbf{V}(\bar{t})=\frac{A_{\mathrm{d}}}{2}e^{(-\xi_{\mathrm{d}}\Omega_{n,\mathrm{d}}+i\Omega_{\mathrm{d}})\bar{t}}\hat{\mathbf{V}}_{\mathrm{d}}+\frac{B_{\mathrm{d}}}{2}e^{i\Omega_{\mathrm{d}}\bar{t}}\hat{\mathbf{V}}_{\mathrm{d}}+\mathrm{c.c.},$ into Eq. (A12):
$\frac{B_{\mathrm{d}}}{2}\left(-\Omega_{\mathrm{d}}^{2}\mathbf{M}+i\Omega_{\mathrm{d}}\mathbf{C}+\mathbf{K}\right)\hat{\mathbf{V}}_{\mathrm{d}}e^{i\Omega_{\mathrm{d}}\bar{t}}+\mathrm{c.c.}=\frac{\mathbf{F}_{0}}{2}e^{i\Omega_{\mathrm{d}}\bar{t}}+\mathrm{c.c.},$ (A15)
$\left[-b_{R}\Omega_{\mathrm{d}}^{2}\mathbf{M}-b_{I}\Omega_{\mathrm{d}}\mathbf{C}+b_{R}\mathbf{K}\right]\hat{\mathbf{V}}_{\mathrm{d}}=\mathrm{Re}(\mathbf{F}_{0});$ (A16a)
$\left[-b_{I}\Omega_{\mathrm{d}}^{2}\mathbf{M}+b_{R}\Omega_{\mathrm{d}}\mathbf{C}+b_{I}\mathbf{K}\right]\hat{\mathbf{V}}_{\mathrm{d}}=\mathrm{Im}(\mathbf{F}_{0}),$ (A16b)
For a sinusoidal forcing ($\mathrm{Re}(\mathbf{F}_{0})=0$), we get:
$a_{R}=-b_{R}=-b_{I}\Omega_{\mathrm{d}}\alpha,$
where, $\alpha=\hat{\mathbf{V}}_{\mathrm{d}}^{\mathrm{T}}\left[\mathbf{K}-\Omega_{\mathrm{d}}^{2}\mathbf{M}\right]^{-1}\mathbf{C}\hat{\mathbf{V}}_{\mathrm{d}}$, and,
$a_{I}=\frac{\xi_{\mathrm{d}}\Omega_{n,\mathrm{d}}a_{R}}{\Omega_{\mathrm{d}}}-b_{I}\Longrightarrow a_{I}=-b_{I}(1+\xi_{\mathrm{d}}\Omega_{n,\mathrm{d}}\alpha),$
$b_{I}=\left[\frac{\alpha}{1+\alpha^{2}\Omega_{\mathrm{d}}^{2}}\right]\hat{\mathbf{V}}_{\mathrm{d}}^{\mathrm{T}}\mathbf{C}^{-1}\mathrm{Im}(\mathbf{F}_{0}),$
Therefore, the analytical form of the DMoA PnC response is given as:
$\mathbf{V}(\bar{t})=-\left(\frac{b_{I}}{2}\right)\left[\alpha\Omega_{\mathrm{d}}+i(1+\xi_{\mathrm{d}}\Omega_{\mathrm{n,d}}\alpha)\right]e^{(-\xi_{\mathrm{d}}\Omega_{\mathrm{n,d}}+i\Omega_{\mathrm{d}})\bar{t}}\hat{\mathbf{V}}_{\mathrm{d}}+\left(\frac{b_{I}}{2}\right)(i+\alpha\Omega_{\mathrm{d}})e^{i\Omega_{\mathrm{d}}\bar{t}}\hat{\mathbf{V}}_{\mathrm{d}}+\mathrm{c.c.},$
that, yields the steady state ($\bar{t}\rightarrow\infty$) response:
$\mathbf{V}(\bar{t})=$ $\left(\frac{b_{I}}{2}\right)(i+\alpha\Omega_{\mathrm{d}})e^{i\Omega_{\mathrm{d}}\bar{t}}\hat{\mathbf{V}}_{\mathrm{d}}+\mathrm{c.c.}$
$=$ $b_{I}\left[\alpha\Omega_{\mathrm{d}}\cos(\Omega_{\mathrm{d}}\bar{t})-\sin(\Omega_{\mathrm{d}}\bar{t})\right]\hat{\mathbf{V}}_{\mathrm{d}}$
$\dot{\mathbf{V}}=$ $-b_{I}\Omega_{\mathrm{d}}\left[\alpha\Omega_{\mathrm{d}}\sin(\Omega_{\mathrm{d}}\bar{t})+\cos(\Omega_{\mathrm{d}}\bar{t})\right]\hat{\mathbf{V}}_{\mathrm{d}}$ (A17)
A5 Multiple defects
@@
Figure A3: 1D Multi-DMoA PnC with defect frequencies in $2^{\mathrm{nd}}$ BG of the MoA PnC. (a) Absolute percentage error contours for a Multi-DMoA PnC ($\mathrm{N}=21$) with two defects at $x_{\mathrm{d1}},x_{\mathrm{d2}}\in[0,\mathrm{N}-1]$ and $x_{\mathrm{d1}}\neq x_{\mathrm{d2}}$, (b) Eigenmode shapes of the Multi-DMoA PnC with one edge ($x_{\mathrm{d1}}=0$) and one bulk ($x_{\mathrm{d2}}$) defect. Multi-DMoA PnC material parameters - $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$, $k_{\mathrm{r}}=1$ and $\{m_{r},k_{\mathrm{gd,r}}\}=\{1,6\}$
@@
The modified analytical approach presented in this article can be applied to determine the defect resonance frequencies of a Multi-DMoA PnC provided the additional defects remain relatively dormant at the resonance frequency of a considered defect. To verify the accuracy of the analytical estimate of $\Omega_{\mathrm{d}}$, Fig. 4 in the main text plots the absolute percentage error b.w. the numerical and analytical values of the defect frequency given by,
@@
When the defect frequencies lie in the $1^{\mathrm{st}}$ BG of the MoA PnC, we notice a small surge in the $\Omega_{\mathrm{d1}}$ error when at least one of the penultimate masses are prescribed defects, i.e., $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$. This error surge is attributed to the spatial in-phase vibration associated with the defect modes in the $1^{\mathrm{st}}$ BG and the significant impact of defect proximity to the finite boundary. Alternately, Fig. A3 explores a Multi-DMoA PnC with defect modes now in the $2^{\mathrm{nd}}$ BG of a MoA PnC. The error contours at $x_{\mathrm{d1,d2}}=1\ \mathrm{or}\ \mathrm{N}-2$ in Fig. A3a are virtually indistinguishable from the surrounding configurations (blue region - $0.003\%$, surge region - $0.29\%$) as opposed to the error contours in Fig. 4a (blue region - $0.004\%$, surge region - $3.9\%$). The reduction in the error surge is indeed a result of improved accuracy of the analytical model and out-of-phase vibration of the defect and boundary masses as the absolute error, $\Delta\Omega_{\mathrm{d}}$ also decreases for defect modes in the $2^{\mathrm{nd}}$ BG.
Figure A4: 3D model of a PnC with a virtual ground.
A6 PnC with a Virtual Ground
Fig. A4 shows a mechanical 3D model with rigid aluminum plates, elastic silicone beams, Veroclear grounding clips, and a rigid (metal) frame surrounding the periodic plate-beam-clip lattice, enabling a virtual grounding connection. The finite resonance frequencies of the PnCs with a virtual ground correspond to the eigenvalues of $\mathbf{K}_{\mathrm{d}}=\mathbf{K}_{\mathrm{VG}}-\Omega^{2}\mathbf{M}$, where the modified stiffness matrix,